OpenStudy (anonymous):

If F(x) = x^2 + 3x then simplify f(x+h) - f(x)/h

Nnesha (nnesha):

If F(x) = x^2 + 3x FIRST YOU HAVE TO REPLACE X BY (X+H)

Nnesha (nnesha):

first you have to replace x by (x+h) sorry caps lock was on btw i'm not yelling

OpenStudy (anonymous):

so then it is (x+h)^2 + 3(x+h)

Nnesha (nnesha):

yes right now solve this

Nnesha (nnesha):

(x+h)^2 same as (x+h)(x+h) so you have to foil it for 3(x+h) you have to distribute parenthesis by 3

OpenStudy (anonymous):

okay give me one second to do it out

Nnesha (nnesha):

np :)

OpenStudy (anonymous):

x^2 + xh + hx+ h^2 + 3x + 3h

Nnesha (nnesha):

good now combine like terms

Nnesha (nnesha):

xh+xh = ??

OpenStudy (anonymous):

they go away because there the same correct?

Nnesha (nnesha):

nope bec both are positive

OpenStudy (anonymous):

so its xh^2

Nnesha (nnesha):

nope when you multply variables then you should add their exponent you are adding xh+xh

Nnesha (nnesha):

how many xh are there ??

OpenStudy (anonymous):

2 so would it be 2xh

Nnesha (nnesha):

yes right

Nnesha (nnesha):

now use that formula and replace f(x+h) by x^2 + 2xh+ h^2 + 3x + 3h and f(x) by x^2 +3x

OpenStudy (anonymous):

I'm sorry what?

Nnesha (nnesha):

okay so f(x+h) = x^2 + 2xh+ h^2 + 3x + 3h and f(x) = x^2 +3x(original question f(x))

OpenStudy (anonymous):

every x replace (x+h)

Nnesha (nnesha):

 use this formula